Gas is confined in a tank at a pressure of 11.0 atm and a temperature of 25.0°C. If two-thirds of the gas is withdrawn and the temperature is raised to 75.0°C, what is the pressure of the gas remaining in the tank?



1
Expert's answer
2022-02-28T10:06:40-0500

Solution

the pressure of the gas remaining in the tank is given by

"P'=\frac{T'VP}{TV'}\\=\frac{75*V*11}{25*V/3}\\=99atm"